OpenStudy (anonymous):

Use the properties of logarithms to write three different expressions equal to log (24). At least one expression should be a sum and one should be a difference.

OpenStudy (apoorvk):

log24= log(6*40)=log6 + log4 log24 =log(48/2)=log48 - log2 log24=-log(1/24)